[ Bob has enough fertilizer to cover 310 ft^2 of garden space. What is the diameter of the largest circular garden that Bob can fertilize? Use 3.14 as pie. ]
Answer by TimothyLamb(4379)

---
circle area:
a = pi*r*r
---
r*r = a/pi
r = sqrt(a/pi)
---
d = 2*r
d = 2*sqrt(a/pi)
d = 2*sqrt(310/3.14)
---
solution:
d = 19.87 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php